OpenStudy (hellomynameiskam):

What is the difference? 3/2 - 7/10 A.- 3/10 B. - 2/5 C. 2/5 D. 4/5 @rebeccaxhawaii and @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

equal the denominators

OpenStudy (alexandervonhumboldt2):

the lest common is 10

OpenStudy (alexandervonhumboldt2):

so we have 15/10 - 7/10=(15-7)/10

OpenStudy (hellomynameiskam):

8/10

OpenStudy (alexandervonhumboldt2):

yes but this can be simplfied

OpenStudy (alexandervonhumboldt2):

\[\frac{ 8 }{ 10 }=\frac{ 4*2 }{ 5*2 }=\frac{ \cancel 2*4 }{ \cancel 2*5 }\]

OpenStudy (hellomynameiskam):

so its D?

OpenStudy (just_one_last_goodbye):

Someone called NightWing? c;'

OpenStudy (just_one_last_goodbye):

@AlexandervonHumboldt2 you got this? ^_^

OpenStudy (alexandervonhumboldt2):

yes D

OpenStudy (anonymous):

D